brandon has 20 coins consisting of nickels and dimes. if the nickels were dimes and the dimes were nickels, he would have 30 cents more than he has now. how many dimes did he have to begin with?
Answer by macston(5194)

.
N=nickels; D=dimes
.
N+D=20
D=20-N
.
$0.05N+$0.10D=$0.10N+$0.05D-$0.30
$0.05D=$0.05N-$0.30
$0.05(20-N)=$0.05N-$0.30
$1.00-$0.05N=$0.05N-$0.30
$1.30=$0.10N
13=N
ANSWER 1: He has 13 nickels.
.
D=20-N=20-13=7
ANSWER 2: He had 7 dimes.
.
CHECK:
.
$0.05N+$0.10D=$0.10N+$0.05D-$0.30
$0.05(13)+$0.10(7)=$0.10(13)+$0.05(7)-$0.30
$0.65+$0.70=$1.30+$0.35-$0.30
$1.35=$1.35